# Ferromagnetic Resonance in a Magnetically Dilute Percolating Ferromagnet: An Experimental and Theoretical Study

### v2

@@ I Introduction
Achieving direct spin control in optoelectronic devices would mark a significant technological breakthrough. In this context dilute ferromagnetic semiconductors (DFS) have attracted substantial research interest because they combine the functionalities of semiconductors (or insulators) with spin degrees of freedom [1, 2], thus paving the way to technological applications [3]. Within the systems being considered [4], (Ga,Mn)N or Ga1-xMnxN, with $x$ representing the fraction of manganese ions distributed randomly on gallium sites, has gained substantial interest. This is mainly due to the high thermal stability of the wide band gap GaN host, that makes it suitable for use in extreme environments, such as outer space or harsh industrial settings.
Despite numerous attempts [5], achieving room-temperature ferromagnetism (FM) in DFS remains an elusive goal in materials science. In the case of Mn-doped GaN, this is primarily due to the insulating character of both GaN:Mn [6] and (Ga,Mn)N mixed crystals [7, 8, 9], which precludes magnetic coupling via itinerant carriers. In the wurtzite GaN lattice Mn is tetrahedrally coordinated by 4 neighboring nitrogen ions. In the absence of other intentional doping or charged point defects, Mn assumes a neutral acceptor oxidation state, Mn3+ ($d^{4}$ configuration) [10, 11], whose dominant coupling mechanism is short-range superexchange [12, 7, 13]. This coupling has been confirmed to be ferromagnetic for at least 25 coordination shells [14].
These unique physical conditions allow the formation of a network in which FM interactions can propagate in a percolation-like manner across the real samples at nonzero temperatures [15, 13, 16, 17, 18], even for Mn concentrations well below 18%, the percolation threshold for the nearest-neighbor-only scenario. We owe it to the fact that although the exchange integral decreases strongly with distance ($r$) across further coordination shells, at the same time the number of possible Mn ions in those shells increases. Overall, the effective coupling in (Ga,Mn)N decays slower with distance, approximately as $\exp(-r/b)$, with $b$ as high as $1.1$  [19]. This situation is very different from the canonical dilute ferromagnetic semiconductor (Ga,Mn)As [20, 21], where a robust global FM state can be stabilized above 170 K [22] by hole-mediated, long-range coupling operating over the same range of Mn compositions [2].
There are a few profound implications of the percolating form of FM in (Ga,Mn)N to our study. Whereas all of the spins are expected to be coupled at $T=0$, only a fraction of the spins present in the material supports the FM state at nonzero temperatures. Starting from high temperatures, Mn spins form clusters of various shape and sizes and the Curie temperature $T_{\mathrm{C}}$, marks that temperature at which the first infinite, percolating ("wall-to-wall") FM cluster forms. Therefore, at nonzero temperatures such a magnetic system is expected to behave as an ensemble of infinite and finite FM clusters (macrospins), characterized by a wide range of blocking temperatures and relaxation times.
This makes the magnetism, in general, and $T_{\mathrm{C}}$, in particular, to depend on the characteristic measurement time ($t_{\mathrm{m}}$) of the experimental technique in use. Similar to the blocking behavior in magnetically granular systems, the longer the $t_{\mathrm{m}}$, the lower the characteristic temperature (in this case $T_{\mathrm{C}}$), as more macrospins are given sufficient amount of time to reach equilibrium, in which they average out their contributions to the measured magnetic properties (e.g. magnetic moment $m$ in case of magnetometry). Consequently, the magnetism—and the determination of $T_{\mathrm{C}}$—depends on $t_{\mathrm{m}}$ due to macrospin-relaxation processes. Interestingly, this phenomenon was also observed in (Ga,Mn)As, where, due to various technological factors [23, 24] or deliberate measurement design [25, 26, 27], the material can be pushed toward the localization limit of holes. At this boundary, magnetic coupling is maintained only on mesoscopic scales [28, 29, 25, 30].
Another important factor is that Mn ions in GaN retain their single-ion magnetic anisotropy (SIMA), which is only weakly modified by their mutual FM coupling. Mn3+ ions in GaN adopt a high-spin configuration ($S=2$, $L=2$), with the ${}^{5}T_{2}$ state being the ground state. In the paramagnetic regime, a strong uniaxial magnetic anisotropy (UMA) emerges at low temperatures, differentiating the magnetic response depending on whether the external magnetic field $\mathbf{H}$ is applied parallel or perpendicular to the hexagonal $c$-axis of the GaN crystal, i.e., $\mathbf{H}\parallel\mathbf{c}$ or $\mathbf{H}\perp\mathbf{c}$, respectively. This anisotropy arises from the trigonal distortion inherent to the wurtzite GaN lattice structure.
In the ideal tetrahedral cooridination the ratio of the hexagonal lattice parameters $c/a=\sqrt{8/3}$. In pure GaN and at low Mn concentrations, $c/a<\sqrt{8/3}$, resulting in UMA with an easy-plane character, favoring spin alignment within the basal plane and making the $c$-axis the hard axis. However, upon increasing Mn concentration, the situation changes significantly. For Mn content, $x_{\mathrm{Mn}}$, exceeding about 10%, a transition to an elongated tetrahedron is anticipated in (Ga,Mn)N [16], which leads to a tensile trigonal distortion capable of reversing the easy magnetization direction to the $c$-axis.
Between these two regimes, an intriguing scenario may arise where the trigonal distortion is marginal, effectively nullifying UMA. Under such conditions, a significantly weaker threefold magnetic anisotropy could be potentially observed, originating from the static Jahn-Teller (JT) effect. This effect displaces N ions surrounding the Mn ion, thereby lowering its site symmetry from trigonal to rhombic [31, 32]. Although the JT-related anisotropy is relatively weak in the ground state, it plays a pivotal role in magnetization switching processes in FM (Ga,Mn)N [33].
Prior studies have definitively shown the importance of investigating static micro-magnetic properties in (Ga,Mn)N layers [34, 17, 13, 7]. The manipulation of magnetic anisotropy in (Ga,Mn)N using electric fields was successfully demonstrated in both the paramagnetic [35] and FM case [33]. Furthermore, using the spin Hall magnetoresistance technique [36], the spectrum of (Ga,Mn)N functionalities was recently supplemented by the demonstration of efficient generation and detection of spin currents in Pt/(Ga,Mn)N structures up to 50 K [37], i.e., well above the Curie temperature established upon magnetometry method. Despite these advancements, our understanding of the dynamics and relaxation processes of magnetization within (Ga,Mn)N, along with the intricate physics governing magnetic anisotropy, remains in its infancy. To address these shortcomings, the ferromagnetic resonance (FMR) technique is employed, a highly precise method renowned for its capability [widely used for its ability] to probe the dynamic magnetic properties of ferromagnetic thin films [38].
FMR is the ideal method for comprehensively assessing fundamental characteristics, including absolute values of magnetic anisotropies, Curie temperature, magnetoelastic coupling coefficients, and parameters describing magnetization relaxation [39]. While FMR techniques have been instrumental in exploring the dynamics and relaxation processes of magnetization in other DFS [39, 40, 41, 42], the realm of (Ga,Mn)N remains largely uninvestigated in this context. This is primarily due to the fact that, unlike Mn2+ ($d^{5}$ configuration), the paramagnetic Mn3+ ion in GaN tetrahedral coordination is EPR silent in typical X-band spectrometers. Consequently much of the information available today on the properties of the Mn3+ state in GaN was established by optical spectroscopy [43, 44] and low-temperature and high-field magnetometry [43, 31, 34, 45]. In this context, the robust FM state observed in (Ga,Mn)N [17] allows us to address the material using classical EPR equipment by studying FMR, the collective response of the otherwise EPR-silent Mn3+ ions. Our study, therefore, aims to expand the pool of understanding of this dilute magnetic percolating system well beyond the limits of static SQUID measurements ($t_{\mathrm{m}}$ of the order of seconds) into the realm of GHz, specific to FMR technique.
@@
There exists a range of well-established and successfully tested theoretical approaches that describe individual aspects of (Ga,Mn)N magnetism. The crystal-field model (CFM), originally developed for Cr2+ in II-VI semiconductors [46, 47, 48] and later adapted to Mn3+ in GaN [49, 43, 44, 31, 34, 35, 50, 33], provides precise predictions of $m(H)$ in high-field conditions but treats Mn ions essentially as non-interacting. Extending this approach to include interactions involves constructing a Hamiltonian matrix whose dimension grows exponentially with the number of Mn spins included, i.e., $25^{n}\times 25^{n}$, where $n$ is the number of nearest-neighbor spins considered in the system. This renders the computation intractable beyond $n\approx 4$, and omits interactions with more distant neighbors entirely [7, 50, 51, 33]. While this method can achieve good agreement with experimental data for fields above about 30 kOe [50], it fails to capture magnetization at low fields. Notably, CFM predicts vanishing remnant magnetization, $M_{\mathrm{REM}}$ = 0, at any temperature, in contradiction with experimental data showing $M_{\mathrm{REM}}\simeq 20$% of $M(H\simeq 50\leavevmode\nobreak\ \mathrm{kOe})$ at $T=2$ K [17, 18] which gradually decreases with temperature and disappears at $T_{\mathrm{C}}$. To account for ferromagnetic ordering and the Curie temperature in (Ga,Mn)N, an alternative approach based on Monte Carlo simulations was adopted. This method employs tight-binding-derived exchange integrals combined with Metropolis spin updates and has been shown to reproduce the experimental $T_{\mathrm{C}}$ values accurately across the experimentally accessible range $0.01\leq x\leq 0.10$ [52, 17], outperforming mean-field approaches based on ab initio inputs, which tend to overestimate $T_{\mathrm{C}}$ by up to an order of magnitude [53]. Nevertheless, neither of these approaches alone is capable of simultaneously describing the full range of magnetic properties—such as magnetization curves, remnant magnetization, coercive fields, Curie temperatures, and, crucially for this study, collective spin dynamics (i.e., FMR resonance fields and linewidths)—within a single, unified set of parameters. To address this issue, we employ an atomistic spin model [54] based on classical spin vectors, combined with the stochastic Landau–Lifshitz–Gilbert (sLLG) formalism [51, 33], which allows us to capture both equilibrium magnetic properties and precessional dynamics. This framework enables us to explore to what extent a unified model can account for both static magnetization and FMR behavior in the magnetically dilute, percolating ferromagnet (Ga,Mn)N.
In this work, the FMR technique and superconducting quantum interference device (SQUID) magnetometry are employed as complementary methods to probe magnetic properties of the insulating ferromagnetic Ga1-xMnxN layer with an Mn concentration of $x=7.9$%. Our primary focus lies in analyzing the temperature dependence of the resonance field and its correlation with the angular orientation of the applied magnetic field. This approach allows us to determine critical parameters such as the Curie temperature, the effective $g$-factor, and magnetic anisotropy constants across varying temperatures. Moreover, our investigation extends to explore the temperature- and angle-dependent linewidth of the FMR signal. This analysis is crucial for revealing insights into material inhomogeneities and relaxation mechanisms in (Ga,Mn)N. Importantly, we also demonstrate that the broad sets of experimental observables—those obtained via SQUID magnetometry and FMR—can be consistently reproduced using a single atomistic spin model with one common set of parameters. Despite the rudimentary representation of the crystalline structure and Mn ion distribution and other simplifications, this unified approach provides remarkably good agreement with experimental data, underscoring the robustness and applicability of the model in describing of such challenging systems as spin-diluted percolating magnets.
@@ II.2 Numerical simulations
Details of the simulation methodology, including the full Hamiltonian, equations of motion, and parameter values, are provided in Appendix B, while additional discussion of the model assumptions and limitations is given in Appendix C.
@@ III.1 Magnetic characterization
The Curie temperature, $T_{\mathrm{C}}$, is the main parameter describing any ferromagnetic material [62], yet its determination has remained a challenge despite numerous advances in this field. The challenge grows in magnetically dilute materials when the critical features become smeared over a wide temperature range. In this study, we employ the thermoremnant magnetization (TRM) measurements as the least cumbersome method of the determination of $T_{\mathrm{C}}$ in DFS, regardless of the range of the FM interaction [17, 63, 18, 64].
@@
Exemplary magnetization curves, $M(H)$, obtained for the same sample at $T=5$ K $<T_{\mathrm{C}}$ are shown in Fig. 1 (b) as solid symbols. Despite being in the FM state, (Ga,Mn)N exhibits a considerable rounded $M(H)$ at low magnetic fields [15, 16, 17, 18, 65, 50, 33]. This behavior arises directly from the significant magnetic dilution and short-range character of the superexhange interaction, both giving rise to the percolative nature of the FM coupling.
On a microscopic level, this effect manifests itself as a broad distribution of ferromagnetically coupled clusters of varying sizes, meaning that at weak magnetic fields, hysteresis involves only a fraction of the spins. At this level of dilution, (Ga,Mn)N does not form a uniform ferromagnetic phase below $T_{\mathrm{C}}$. Moreover, the $M(H)$ data indicate that magnetization does not fully saturate even at fields as high as 70 kOe. This behavior arises from the significant contribution of the weakly saturating orbital moment to the total magnetic moment of Mn3+ ions [66].
The results of the atomistic spin simulations corresponding to the relevant magnetization data—namely, the $M_{\mathrm{TRM}}(T)$ curve presented in Fig.1(a) and the $M(H)$ isotherms in Fig.1(b)—are shown as open symbols. The simulation methodology used to model both dependencies is described in Appendix B. Importantly, these two qualitatively distinct spin-dynamics regimes are simulated using the same set of parameters, which were initially optimized to reproduce the experimental $M(H)$ curves as accurately as possible. As expected, the agreement between the model and experiment is very good for $M(H)$, particularly in the high-field regime. The main quantitative discrepancies appear in the low-field region (as shown in the inset of Fig. 1(b)) and in the simulation of the $M_{\mathrm{TRM}}(T)$ dependence. In both cases, the model tends to overestimate the magnetization at low fields.
These discrepancies originate from three principal limitations of the current modeling approach, which are outlined in more detail in Appendix C. Specifically: (i) the model does not account for pairwise Mn–Mn anisotropy effects [32]; (ii) Jahn–Teller-type anisotropy is approximated using a cubic term; (iii) distribution of Mn ions is assumed to be perfectly random. All these factors contribute to a modified anisotropy landscape, which becomes particularly relevant at low magnetic fields, where the Zeeman energy is insufficient to fully suppress local anisotropy effects. Consequently, the simulated magnetization tends to deviate more strongly from experiment in this regime.
Consequently, achieving a Curie temperature differing from experiment by only a factor of two can be regarded as a reasonably good result. We mention that doubling the magnitude of exchange integral would indeed reproduce the experimental $T_{\mathrm{C}}$, but this would produce almost rectangular hysteresis loops with greatly overestimated coercive fields—an outcome clearly at odds with the experiment.
Overall, even though our computational approach captures only a rudimentary level of the real physical complexity in (Ga,Mn)N, the agreement achieved here remains quite notable. It therefore serves as a promising basis for modeling the FMR response of (Ga,Mn)N, as presented in Secs. III.3 and III.4.
@@ III.2 Temperature dependence of the FMR
Fig. 2 (a) depicts the temperature dependence of the ferromagnetic resonance spectra when $\mathbf{H}$ is applied along the out-of-plane direction ($\mathbf{H}\parallel[0001]$). The hard direction was chosen to minimize the effects of superposition of the FMR signal with the cyclotron resonance (CR) background signal stemming from the conducting GaN template. In contrast to SQUID results, at low temperatures (up to 7 K) only a strong CR and a very weak defect-related paramagnetic signal at $g=2$ are visible. Ferromagnetic signatures start to appear around $T=9$ K and their integrated intensity gradually decreases with increasing temperature to disappear entirely above $T=70$ K. This behavior is markedly different from that observed in other DFSs, such as (Ga,Mn)As [39, 67, 41], (Ga,Mn)P [42], or the (Pb,Mn)(Te,Se) family [68], where well-defined FMR signals are routinely detected far below the Curie temperature. We attribute the absence of detectable FMR signals below $T_{\mathrm{C}}$ in our (Ga,Mn)N sample primarily to strong homogeneous broadening, discussed in detail in Section III.4. The underlying mechanism remains to be fully elucidated, but it is likely related to an enhanced relaxation rate of the magnetization that causes the resonance to broaden beyond detection limit. A plausible origin of this rapid relaxation lies in the high density of structural defects—particularly threading dislocations—commonly present in (Ga,Mn)N films grown on foreign substrates like sapphire [69, 65].
In order to compare the FMR results with the magnetization measured with SQUID at nonzero magnetic fields we fit the signals with the derivative of a Lorentzian function:
@@
to establish the amplitude $A$, resonance field $H_{R}$, and the half width at half maximum (HWHM) $\Delta H$ at each temperature. In the fits we also include the influence of the substrate free electron absorption, which at high enough fields can be approximated by the sum of a decreasing exponential function and a linear increase with $H$. The fits are denoted by black dashed lines in Fig. 2 (a). The FMR intensity, $I$, is then calculated as $I=A\times\Delta H$ and is directly proportional to the magnitude of the ferromagnetic moment. In the absence of magnetocrystalline anisotropy it would be possible to obtain absolute values of magnetization from the in- and out-of-plane FMR measurements. This is not the case here. Therefore, we adjust the established intensity of the FMR signal to the magnetization value determined from SQUID at 3 kOe and 55 K. As shown in Fig. 2 (b) a close correspondence is observed between the $M(T)$ values measured between $H=3$ and 4 kOe and $I(T)$ (pentagons) derived from the FMR data. In both cases the transition temperature exceeds $T_{\mathrm{C}}$. The fact that the FMR signal persists up to $T=70$ K is attributed to the magnetic field ($H=3.3$ kOe) applied in the FMR experiment, which acts against thermal fluctuations [40, 70], and the presence of non-percolating ferromagnetic clusters above $T_{C}$.
@@
To substantiate this scenario we re-draw the magnetometry and FMR results in a double logarithmic scale, shown in Fig. 3, since in amorphous systems with a wide range of exchange integrals the magnetic susceptibility, $\chi(T)=M(T)/H$, above their specific coupling temperature is expected to be proportional to $T^{-\alpha}$ [71, 72]. Obviously, $\alpha=1$ for non-interacting systems (the Curie law) and this dependency is marked by a dashed line in Fig. 3. However, the presence of interactions tends to form spin pairs and higher order complexes which reduce/enhance the magnetization on cooling with respect to $T^{-1}$ for antiferromagnetic/ferromagnetic coupling among the spins. As a result $\alpha<1$ for random antiferromagnets. Here Mn- or Co-based II-VI dilute magnetic semiconductors are good examples [73, 74, 75]. Analogously, $\alpha>1$ for ferromagnetic coupling, and (Ga,Mn)N is the prime example here [17]. Our magnetometry data plotted in Fig. 3 follow the same trend as in [17] yielding $\alpha=1.23$ for $T>80$ K (marked by the solid black line). This behavior is obeyed at high temperatures independently of the magnitude of $H$, at least within 1 to 10 kOe range.
However, below $T_{f}=70$ K, the entire set of curves deviates upward from the $T^{-1.23}$ trend and begins to fan out below 30 K, without exhibiting any signatures of critical behavior near $T_{\mathrm{C}}$. The FMR intensity follows a similar trend, with the important observation that the temperature at which the FMR signal rises above the noise level coincides with $T_{f}$, i.e. the same temperature at which $\chi(T)$ begins to exceed the values predicted by the quasi-paramagnetic $T^{-1.23}$ dependence. This non-coincidental agreement suggests that $T_{f}$ marks the onset of formation of sufficiently large ferromagnetically ordered clusters—macrospins—that coherently absorb microwave energy.
@@
On the other hand, Mn3+ ions in GaN have a tendency to cluster [76], and minimizing this effect to achieve a homogeneous Mn distribution required several years of intensive materials research [15, 77, 16, 18]. Despite this substantial technological progress, the resulting layers still exhibit measurable macroscopic inhomogeneities in the spin distribution, yet characterized by a rather small variance $\Delta x\simeq 0.2$% [17]. This value appears too low to reflect the presence of relatively small regions with strongly enhanced Mn concentration, as would be expected for spinodal decomposition (chemical phase separation) [78], which is ubiquitously responsible for the room-temperature ferromagnetism in (Ga,Mn)N and related compounds. Resolving this issue requires probes with atomic resolution, such as Atom Probe Tomography, and thus lies beyond the scope of this work.

### v3

@@ I Introduction
Achieving direct spin control in optoelectronic devices would mark a significant technological breakthrough. In this context, dilute ferromagnetic semiconductors (DFS) have garnered considerable attention due to their unique combination of semiconductor functionality with spin degrees of freedom [1, 2], opening avenues for spintronic applications [3]. Among these, Ga1-xMnxN has attracted particular interest due to the robustness of the wide-bandgap GaN host, which makes it suitable for harsh environments.
Despite numerous efforts [4], achieving room-temperature ferromagnetism (FM) in DFS remains a major challenge. In Mn-doped GaN, this difficulty arises from the insulating character of both GaN:Mn [5] and (Ga,Mn)N [6, 7, 8], which precludes carrier-mediated magnetic coupling. In the wurtzite GaN lattice, Mn3+ ions ($d^{4}$ configuration) couple primarily via short-range ferromagnetic superexchange [9, 6, 10], with interactions persisting over many coordination shells [11]. This long-range nature, combined with the geometric increase in Mn neighbors at greater distances, results in an effective coupling that decays approximately as $\exp(-r/b)$ with $b\approx 1.1$  [12]. Consequently, a percolative ferromagnetic state emerges even at low Mn concentrations [13, 10, 14, 15, 16], in contrast to the hole-mediated long-range FM of (Ga,Mn)As [17, 18, 19].
The percolative nature of FM in (Ga,Mn)N gives rise to a heterogeneous ensemble of magnetic clusters, both finite and infinite (macrospins), whose contributions depend on temperature and the characteristic measurement time. Interestingly, this phenomenon was also observed in (Ga,Mn)As, where, due to various technological factors [20, 21] or deliberate measurement design [22, 23, 24], the material can be pushed toward the localization limit of holes. This specific dynamic behavior [25, 26, 27] directly impacts quantities such as Curie temperature ($T_{\mathrm{C}}$), which become protocol-dependent.
Mn3+ ions in GaN also exhibit strong single-ion magnetic anisotropy, stemming from trigonal crystal field distortions intrinsic to the wurtzite structure [28, 29]. At low Mn concentrations, this anisotropy favors easy-plane magnetization. Moreover, Jahn–Teller effects may introduce additional threefold in-plane anisotropy, which is particularly relevant for magnetization switching in the ferromagnetic state [30].
Previous studies have established key static magnetic properties of (Ga,Mn)N [31, 15] and recent work has demonstrated the electric field control of magnetic anisotropy [32, 30] and the presence of spin Hall magnetoresistance in Pt/(Ga,Mn)N heterostructures [33]. Yet, dynamic properties such as spin relaxation and magnetization precession remain underexplored. Ferromagnetic resonance (FMR) is a powerful tool to probe these dynamics [34, 35], but has rarely been applied to (Ga,Mn)N due to the EPR silence of Mn3+ ions in typical X-band spectrometers [36, 37]. However, the presence of a robust ferromagnetic state allows for FMR detection of collective excitations, offering a route to study dynamic behavior in this otherwise EPR-silent system [15, 38].
@@
On the theoretical ground, there exists a range of well-established and successfully tested approaches that describe individual aspects of (Ga,Mn)N magnetism. Full quantum mechanical approaches such as the crystal-field model (CFM) - originally developed for Cr2+ in II-VI semiconductors [39, 40, 41] and later adapted to Mn3+ in GaN [42, 36, 37, 28, 31, 32, 43, 30], provide accurate descriptions of high-field magnetization but treats Mn ions essentially as non-interacting. Extending this approach to include interactions involves constructing a Hamiltonian matrix whose dimension grows exponentially with the number of Mn spins included, i.e., $25^{n}\times 25^{n}$, where $n$ is the number of nearest-neighbor spins considered in the system. This renders the computation intractable beyond $n\approx 4$, and omits interactions with more distant neighbors entirely [6, 43, 44, 30]. While this method can achieve good agreement with experimental data for fields above about 30 kOe [43], it fails to capture magnetization at low fields. Notably, CFM predicts vanishing remnant magnetization, $M_{\mathrm{REM}}$ = 0, at any temperature, in contradiction with experimental data showing $M_{\mathrm{REM}}\simeq 20$% of $M(H\simeq 50\leavevmode\nobreak\ \mathrm{kOe})$ at $T=2$ K [15, 16] which gradually decreases with temperature and disappears at $T_{\mathrm{C}}$. Monte Carlo simulations offer insight into $T_{\mathrm{C}}$ [45], outperforming mean-field approaches based on ab initio inputs which tend to overestimate $T_{\mathrm{C}}$ by up to an order of magnitude [46], but do not capture FMR dynamics. To address this, we employ a classical atomistic spin model based on the stochastic Landau–Lifshitz–Gilbert (sLLG) formalism [47, 44, 30], which allows us to capture both equilibrium magnetic properties and precessional dynamics, thus paving the way toward a unified simulation of both static and dynamic responses.
In this work, we combine FMR and superconducting quantum interference device (SQUID) magnetometry to investigate a Ga1-xMnxN layer with $x=7.9$%. We analyze the temperature and angular dependence of the resonance field and linewidth, extract anisotropy constants, and evaluate spin relaxation mechanisms. Importantly, we demonstrate that both FMR and SQUID data can be consistently described using a single set of parameters within our atomistic spin model, thus bridging the gap between static and dynamic magnetic characterizations of this spin-diluted percolating system.
This paper is organized as follows. In Section II, we describe the experimental procedures, sample preparation, and introduce the atomistic spin simulations used to model both static and dynamic magnetic properties of (Ga,Mn)N. Section III presents our key findings, starting from the static magnetization data and Curie temperature analysis, followed by a detailed temperature- and angle-resolved FMR study. The section also includes a discussion of magnetocrystalline anisotropy, spin relaxation mechanisms, and a comparison between experiment and sLLG simulations. Finally, Section IV summarizes the main conclusions of the study.
@@ II.2 Numerical simulations
Details of the simulation methodology, including the full Hamiltonian, equations of motion, parameter values, and numerical integration scheme are provided in Appendices B and C, while additional discussion of the model assumptions and limitations is given in Appendix D.
@@ III.1 Magnetic characterization
The Curie temperature, $T_{\mathrm{C}}$, is a key parameter of ferromagnetic materials [55], but its determination remains challenging—especially in magnetically dilute systems, where critical features are often smeared over a broad temperature range. In this study, we employ thermoremnant magnetization (TRM) measurements as a robust method for determining $T_{\mathrm{C}}$ in dilute ferromagnetic semiconductors, regardless of the FM interaction range [15, 56, 16, 57].
@@
Exemplary magnetization curves, $M(H)$, obtained for the same sample at $T=5$ K $<T_{\mathrm{C}}$ are shown in Fig. 1 (b) as solid symbols. Despite being in the FM state, (Ga,Mn)N exhibits markedly rounded $M(H)$ curves at low fields [13, 14, 15, 16, 58, 43, 30]. This behavior stems from magnetic dilution and the short-range nature of superexchange interactions.
On a microscopic level, this leads to a broad distribution of ferromagnetically coupled clusters, where hysteresis at low fields involves only a fraction of spins. At this dilution level, (Ga,Mn)N does not form a uniform FM phase below $T_{\mathrm{C}}$. Moreover, $M(H)$ does not fully saturate even at 70 kOe, reflecting a significant contribution of the weakly saturating orbital moment of Mn3+ ions [59].
The atomistic spin simulations corresponding to the $M_{\mathrm{TRM}}(T)$ and $M(H)$ data in Fig. 1 are shown as open symbols. Both regimes are simulated using a single parameter set, initially optimized to match the experimental $M(H)$ curves. As expected, agreement is very good in the high-field region, while discrepancies arise at low fields (see inset of Fig. 1 (b)) and in the $M_{\mathrm{TRM}}(T)$ simulation. In both cases, the model slightly overestimates magnetization at low fields.
These deviations stem from three limitations of the model, discussed in Appendix D: (i) omission of pairwise Mn–Mn anisotropy effects [29]; (ii) approximation of Jahn–Teller anisotropy with a cubic term; and (iii) assumption of perfectly random Mn distribution. All these factors affect the local anisotropy landscape, especially relevant when Zeeman energy is insufficient to overcome anisotropy barriers.
Achieving a Curie temperature within a factor of two of experiment is thus a reasonable result. We note that doubling the exchange integral would reproduce the experimental $T_{\mathrm{C}}$, but also lead to unrealistic rectangular $M(H)$ loops and overestimated coercive fields — an outcome clearly at odds with the experiment.
Despite its simplicity, the model captures key experimental features and provides a sound basis for simulating the FMR response, as discussed in Sections III.3 and III.4.
@@ III.2 Temperature dependence of the FMR
Fig. 2 (a) shows the temperature dependence of the FMR spectra for $\mathbf{H}\parallel[0001]$. This hard-axis orientation minimizes the overlap between the FMR signal and the cyclotron resonance (CR) background from the conducting GaN template. At low temperatures ($T<7$ K), only a strong CR and a weak $g=2$ paramagnetic signal are detected. Ferromagnetic features emerge near $T=9$ K and persist up to 70 K, with their integrated intensity decreasing with increasing temperature. This behavior contrasts sharply with that observed in other DFSs, such as (Ga,Mn)As [35, 60, 61], (Ga,Mn)P [62], and (Pb,Mn)(Te,Se) [63], where FMR signals are readily observed well below $T_{\mathrm{C}}$.
We attribute the absence of detectable FMR signals below $T_{\mathrm{C}}$ to strong homogeneous broadening, likely caused by an enhanced magnetization relaxation rate, discussed in Section III.4. A plausible origin of this rapid relaxation is the high density of structural defects, particularly threading dislocations, typical for (Ga,Mn)N films grown on foreign substrates [64, 58].
To compare the FMR and SQUID results, we fit the spectra using the derivative of a Lorentzian function
@@
to extract the amplitude $A$, resonance field $H_{R}$, and the half width at half maximum (HWHM) $\Delta H$ at each temperature. The fits are denoted by black dashed lines in Fig. 2 (a). The FMR intensity, $I$, is then calculated as $I=A\times\Delta H$ and is directly proportional to the magnitude of the ferromagnetic moment. While absolute magnetization values cannot be determined due to magnetocrystalline anisotropy, we normalize $I(T)$ to the SQUID magnetization at 3 kOe and 55 K. As shown in Fig. 2 (b), $I(T)$ closely follows $M(T)$ between 3–4 kOe. The persistence of the FMR signal up to 70 K likely reflects both the stabilizing effect of the applied field [65, 66] and the presence of non-percolating FM clusters above $T_{\mathrm{C}}$.
@@
To explore this further, we replot the FMR and SQUID data in a double-logarithmic scale [Fig. 3]. In disordered systems with varying exchange interactions, magnetic susceptibility $\chi(T)=M(T)/H$ follows a power-law: $\chi(T)\propto T^{-\alpha}$ [67, 68]. For noninteracting moments, $\alpha=1$ (Curie law). Antiferromagnetic interactions yield $\alpha<1$ [69, 70, 71], while ferromagnetic ones give $\alpha>1$, as observed in (Ga,Mn)N [15]. Our data follow $\alpha=1.23$ for $T>80$ K, independent of $H$ in the 1–10 kOe range (indicated by the solid black line in Fig. 3).
Below $T_{f}=70$ K, $\chi(T)$ curves deviate upward from the $T^{-1.23}$ trend and begin to fan out below 30 K, without clear critical behavior near $T_{\mathrm{C}}$. Notably, the FMR signal also emerges at $T_{f}$, where $\chi(T)$ begins to exceed the expected quasi-paramagnetic $T^{-1.23}$ response. This coincidence suggests that $T_{f}$ marks the onset of coherent microwave absorption by FM-coupled macrospins.
@@
Although Mn3+ ions in GaN tend to cluster [72], years of materials optimization [13, 73, 14, 16] have led to nearly homogeneous Mn distribution. Residual macroscopic inhomogeneities exhibit a narrow variance $\Delta x\simeq 0.2$% [15], insufficient to indicate phase separation, such as spinodal decomposition [74], known to cause spurious ferromagnetism in related systems. Resolving the fine structure of Mn clustering would require atomically resolved techniques like Atom Probe Tomography, which lie beyond the scope of this work.